OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5).

OpenStudy (ddcamp):

(f°g)(x) is a way of saying f(g(x)). Find g(-5), then find f of your answer to that.

OpenStudy (anonymous):

Do i multiply -2x-7*-4x +3? or do I set it up as -2x-7=-4x+3?

OpenStudy (ddcamp):

Start by taking g of -5: g(-5) = -4(-5) + 3 = 20 + 3 = 23

OpenStudy (ddcamp):

Then take f of that: f(23) = -2(23) - 7 f(23) = -46 - 7 f(23) = -53 So, (f°g)(-5) = f(g(-5)) = f(23) = -53

OpenStudy (anonymous):

Thank you!